OpenStudy (anonymous):

there are 4 grams of fiber in 1/2 cup of oats how many grams of fiber are in 3 1/2 cups of oats A 7 B 12 C 14 D 28

Mehek (mehek14):

first, find out how many cups equal to 1 gram

Mehek (mehek14):

do you know how to do this?

OpenStudy (texaschic101):

4/(1/2) = x/ (3 1/2) cross multiply (turn mixed number to improper fraction ---> 3 1/2 = 7/2 (1/2)x = 4 * 7/2 (1/2)x = 14 -- divide both sides by 1/2 x = 14 / (1/2) x = 14 * 2/1 x = 28 so there is 28 grams of fiber in 3 1/2 cups of oats